# Unifying Deductive and Abductive Reasoning in Knowledge Graphs with Masked Diffusion Model

### v1

@@ Abstract.
Deductive and abductive reasoning are two critical paradigms for analyzing knowledge graphs, enabling applications from financial query answering to scientific discovery. Deductive reasoning on knowledge graphs usually involves retrieving entities that satisfy a complex logical query, while abductive reasoning generates plausible logical hypotheses from observations. Despite their clear synergistic potential, where deduction can validate hypotheses and abduction can uncover deeper logical patterns, existing methods address them in isolation. To bridge this gap, we propose DARK, a unified framework for Deductive and Abductive Reasoning in Knowledge graphs. As a masked diffusion model capable of capturing the bidirectional relationship between queries and conclusions, DARK has two key innovations. First, to better leverage deduction for hypothesis refinement during abductive reasoning, we introduce a self-reflective denoising process that iteratively generates and validates candidate hypotheses against the observed conclusion. Second, to discover richer logical associations, we propose a logic-exploration reinforcement learning approach that simultaneously masks queries and conclusions, enabling the model to explore novel reasoning compositions. Extensive experiments on multiple benchmark knowledge graphs show that DARK achieves state-of-the-art performance on both deductive and abductive reasoning tasks, demonstrating the significant benefits of our unified approach.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 2. Related Work
Abductive Reasoning. Abductive reasoning has been increasingly studied across both natural language and neuro-symbolic domains. In natural language inference, $\alpha$-NLI (Bhagavatula et al., 2020) first introduced abductive reasoning into commonsense reasoning, where plausible explanations are inferred from given observations. Subsequent research has extended this paradigm to handle rare events and logically complex scenarios (Qin et al., 2021; Kadiķis et al., 2022; Chan et al., 2023; Zhao et al., 2024). Benchmarks such as ProofWriter (Tafjord et al., 2021) further evaluate formal abductive reasoning within semi-structured texts containing explicit logical relationships. More recent studies explore the capabilities of large language models in open-world and abstract reasoning settings (Zhong et al., 2023; Del and Fishel, 2023; Thagard, 2024; Liu et al., 2024b; Zheng et al., 2025a). In the neuro-symbolic domain, Abductive Learning (ABL) (Zhou, 2019) bridges machine learning with logical reasoning by integrating perception and symbolic inference. Building upon this foundation, advanced frameworks such as ARLC (Camposampiero et al., 2024) and ABL-Refl (Hu et al., 2025) enhance context-awareness, error correction, and generalization, leading to more efficient and interpretable reasoning processes.
Abductive Reasoning in Knowledge Graphs. The abductive reasoning task in knowledge graphs is defined as generating the most plausible first-order logical hypotheses given a set of observed entities. AbductiveKGR (Bai et al., 2024b) addresses this task using an autoregressive generative paradigm to construct logical hypotheses sequentially. CtrlHGen (Gao et al., 2025), in contrast, introduces a controllable hypothesis generation framework to mitigate the proliferation of redundant or irrelevant hypotheses, enabling the generation of hypotheses that align with user-specified interests. Some studies (Galárraga et al., 2015; Xu et al., 2022; Chen et al., 2022b; Ho et al., 2018) on rule mining aim to discover logical rules within knowledge graphs, a process that can be viewed as a form of inductive reasoning. Insights from these rule mining approaches can also inform and guide abductive reasoning tasks.
@@ 3. Preliminaries
We represent a knowledge graph as $G=(V,R)$, where $V$ denotes the set of entities and $R$ denotes the set of relation types. Each relation $r\in R$ is defined as a binary predicate $r:V\times V\to\{\text{true},\text{false}\}$, which determines whether a directed edge labeled with relation $r$ exists from entity $u$ to entity $v$. In other words, if $r(u,v)=\text{true}$, then the triple $(u,r,v)$ is included in $G$; otherwise, it is not.
Following the open-world assumption of knowledge graphs (Drummond and Shearer, 2006), our methodology is built on the fundamental premise that the observed knowledge graph $G$ available during training is inherently incomplete. Consequently, all unobserved edges or missing triples are treated not as false negatives, but as potentially true facts that are merely unknown. It is formally guaranteed that the observed graph $G$ is a proper subgraph of the complete, underlying knowledge graph $\bar{G}$, which remains hidden and represents the full set of true facts.
@@ 4. Method
In this section, we propose DARK, a masked diffusion model for unifying deductive and abductive reasoning in knowledge graphs. The framework has been shown in Fig 2. In Section 4.1, we first present the core insights that motivate our approach to unifying these two reasoning paradigms. In Section 4.2, we describe how the masked diffusion model serves as the foundation for this unification and propose a self-reflective sampling strategy, where deduction is employed to refine hypotheses in the abductive reasoning process. In Section 4.3, we propose a logic-exploration reinforcement learning method that enables the model to autonomously explore a broader range of logical possibilities, thereby deepening its intrinsic understanding of logic. Finally, in Section 4.4, we analyze the time and space complexity of the proposed self-reflective sampling method.
@@ 4.3. Logic-exploration Reinforcement Learning
4.4. Complexity Analysis
Algorithm 1 Self-Reflective Sampling
Input: Masked state $X^{T}=[\mathbf{M},X_{O}]$, with the known observation $X_{O}$, model $f_{\theta}$, candidate number $p$, reflective steps $T$, mask token $\mathbf{M}$, self-reflective interval $k$.
Output: The generated result $\tilde{X}^{0}$
for $t=T,\dots 1$ do
  if $t\%k=0$ then
   Generate $p$ candidate hypotheses $\{\hat{X}^{(0,i)}\}_{i=1}^{p}$ from $X^{t}$,
   Compute conclusions $\hat{X}_{O}^{(0,i)}=f_{\theta}(\hat{X}_{Q}^{(0,i)},\mathbf{M})$ and similarity scores $s_{i}$,
   Select best query $\tilde{X}_{Q}^{0}=\arg\max_{i}s_{i}$,
   Update the $\widetilde{X}^{0}$ in Eq. (6) and calculate $X^{t-1}$.
  else
   Calculate $\widetilde{X}^{0}=f_{\theta}(X^{t})$,
   Update $X^{t-1}$ directly from Eq. (6).
  end if
end for
return $\tilde{X}^{0}$
@@ 5.3. Deductive Reasoning
Baselines and Metrics.For deductive reasoning, we adopt the complex query answering (CQA) task as our evaluation benchmark. We compare our model with several representative state-of-the-art approaches: BetaE (Ren and Leskovec, 2020), a probabilistic embedding model that uses Beta distributions to represent entities and queries, supporting logical operators while modeling uncertainty. FuzzyQE (Chen et al., 2022a), a query embedding framework based on fuzzy logic, which defines logical operators in a principled, learning-free manner for answering first-order logic queries. Query2Particle (Bai et al., 2022), which encodes each query as a set of particle embeddings, enabling reasoning over arbitrary first-order logic queries by retrieving answers from diverse regions of the embedding space. UlTRAQUERY (Galkin et al., 2024), a foundation model based on GNNs for CQA that captures the four fundamental logical relationships, thereby enabling zero-shot query answering across knowledge graphs. To assess performance on CQA tasks, we employ standard ranking-based metrics: Mean Reciprocal Rank (MRR), which captures the average reciprocal rank of the correct entity, reflecting how highly the model ranks the true answer; and Hits@$k$, which measures the proportion of queries where the correct answer appears in the top-$k$ predictions, indicating the model’s effectiveness in retrieving correct entities among its most confident outputs. Here, since our model cannot directly assign a probability to each entity, we set the probability of the entity answered by the model to 1 and the probability of the others that were not answered to 0. The results have been reported in Table 2.
@@ 6. Conclusion
In this work, we presented DARK, a masked diffusion model to unify deductive and abductive reasoning over knowledge graphs. By modeling bidirectional reasoning as a masked diffusion process, DARK seamlessly integrates deduction and abduction in a mutually reinforcing and dynamically interactive manner, allowing two reasoning directions to refine each other during the iterative denoising process. Our self-reflective denoising mechanism enables iterative hypothesis refinement through deductive validation, while the logic-exploration reinforcement learning strategy further encourages the discovery of richer and more diverse reasoning patterns by jointly masking queries and conclusions. Extensive experiments demonstrate that DARK not only achieves impressive results on multiple benchmark datasets but also exhibits strong generalization ability across diverse reasoning scenarios. This approach leverages the strengths of diffusion models to enhance logical reasoning, offering a robust framework for complex knowledge graph tasks. These findings highlight the potential of integrating diffusion-based modeling with logical reasoning, paving the way for more unified, interpretable, and scalable reasoning frameworks in future research.

### v2

@@ Abstract.
Deductive and abductive reasoning are two critical paradigms for analyzing knowledge graphs, enabling applications from financial query answering to scientific discovery. Deductive reasoning on knowledge graphs usually involves retrieving entities that satisfy a complex logical query, while abductive reasoning generates plausible logical hypotheses from observations. Despite their clear synergistic potential, where deduction can validate hypotheses and abduction can uncover deeper logical patterns, existing methods address them in isolation. To bridge this gap, we propose DARK, a unified framework for Deductive and Abductive Reasoning in Knowledge graphs. As a masked diffusion model capable of capturing the bidirectional relationship between queries and conclusions, DARK has two key innovations. First, to better leverage deduction for hypothesis refinement during abductive reasoning, we introduce a self-reflective denoising process that iteratively generates and validates candidate hypotheses against the observed conclusion. Second, to discover richer logical associations, we propose a logic-exploration reinforcement learning approach that simultaneously masks queries and conclusions, enabling the model to explore novel reasoning compositions. Extensive experiments on multiple benchmark knowledge graphs show that DARK achieves competitive performance on both deductive and abductive reasoning tasks, demonstrating the significant benefits of our unified approach.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792133isbn: 979-8-4007-2307-0/2026/04ccs: Mathematics of computing Graph algorithmsccs: Information systems Information retrieval
@@ 2. Related Work
Abductive Reasoning. Abductive reasoning has been widely studied in both natural language and neuro-symbolic domains. In natural language inference, $\alpha$-NLI (Bhagavatula et al., 2020) introduces abductive reasoning into commonsense inference by generating plausible explanations for given observations, with subsequent work extending this paradigm to rare events and more complex logical settings (Qin et al., 2021; kadiķis2022embarrassinglysimpleperformanceprediction; Chan et al., 2023; Zhao et al., 2024). Benchmarks such as ProofWriter (Tafjord et al., 2021) evaluate formal abductive reasoning over semi-structured texts with explicit logical relations, while recent studies further explore abductive capabilities of large language models in open-world and abstract reasoning scenarios (Zhong et al., 2023; Del and Fishel, 2023; Thagard, 2024; Liu et al., 2024b; Zheng et al., 2025a). In parallel, neuro-symbolic approaches such as Abductive Learning (ABL)(Zhou, 2019) integrate perception and symbolic inference, with extensions like ARLC(Camposampiero et al., 2024) and ABL-Refl (Hu et al., 2025) improving context-awareness, error correction, and generalization. Within knowledge graphs, abductive reasoning focuses on generating plausible first-order logical hypotheses given observed entities. Representative methods include AbductiveKGR (Bai et al., 2024b), which adopts an autoregressive hypothesis generation paradigm, and CtrlHGen (Gao et al., 2025), which introduces controllability to reduce redundant hypotheses. Related studies on rule mining (Galárraga et al., 2015; Xu et al., 2022; Chen et al., 2022a; Ho et al., 2018), although primarily inductive, also provide useful insights for guiding abductive reasoning over knowledge graphs.
@@ 3. Preliminaries
We represent a knowledge graph as $G=(V,R)$, where $V$ is the set of entities and $R$ is the set of relation types. Each relation $r\in R$ is a binary predicate $r(u,v)\in\{\text{true},\text{false}\}$, indicating whether the triple $(u,r,v)$ exists in $G$. Under the open-world assumption (Drummond and Shearer, 2006), the observed graph $G$ is incomplete and contains only a subset of true facts. Unobserved triples are treated as unknown rather than false, and $G$ is assumed to be a subgraph of the complete but hidden knowledge graph $\bar{G}$.
@@ 4. Method
In this section, we propose DARK, a masked diffusion model for unifying deductive and abductive reasoning in knowledge graphs. The framework has been shown in Fig 2. In Section 4.1, we first present the core insights that motivate our approach to unifying these two reasoning paradigms. In Section 4.2, we describe how the masked diffusion model serves as the foundation for this unification and propose a self-reflective sampling strategy, where deduction is employed to refine hypotheses in the abductive reasoning process. In Section 4.3, we propose a logic-exploration reinforcement learning method that enables the model to autonomously explore a broader range of logical possibilities, thereby deepening its intrinsic understanding of logic. Finally, we analyze the time and space complexity of the proposed self-reflective sampling method in Appendix B.
@@ 5.3. Deductive Reasoning
Baselines and Metrics. For deductive reasoning, we adopt the complex query answering (CQA) task as our evaluation benchmark. We compare our model with several representative state-of-the-art approaches: BetaE (Ren and Leskovec, 2020), a probabilistic embedding model that uses Beta distributions to represent entities and queries, supporting logical operators while modeling uncertainty. FuzzyQE (Chen et al., 2022b), a query embedding framework based on fuzzy logic, which defines logical operators in a principled, learning-free manner for answering first-order logic queries. Query2Particle (Bai et al., 2022), which encodes each query as a set of particle embeddings, enabling reasoning over arbitrary first-order logic queries by retrieving answers from diverse regions of the embedding space. UlTRAQUERY (Galkin et al., 2024), a foundation model based on GNNs for CQA that captures the four fundamental logical relationships, thereby enabling zero-shot query answering across knowledge graphs. To assess performance on CQA tasks, we employ standard ranking-based metrics: Mean Reciprocal Rank (MRR), which captures the average reciprocal rank of the correct entity, reflecting how highly the model ranks the true answer; and Hits@$k$, which measures the proportion of queries where the correct answer appears in the top-$k$ predictions, indicating the model’s effectiveness in retrieving correct entities among its most confident outputs. Here, since our model cannot directly assign a probability to each entity, we set the probability of the entity answered by the model to 1 and the probability of the others that were not answered to 0. The results have been reported in Table 2.
@@ 6. Conclusion
In this work, we propose DARK, a masked diffusion model that unifies deductive and abductive reasoning over knowledge graphs. By formulating bidirectional reasoning as a masked diffusion process, DARK enables deduction and abduction to interact and mutually refine each other during iterative denoising. The self-reflective denoising mechanism supports iterative hypothesis refinement via deductive validation, while a logic-exploration reinforcement learning strategy promotes more diverse reasoning patterns through joint masking of queries and conclusions. Extensive experiments on multiple benchmarks demonstrate that DARK achieves strong performance and generalization across diverse reasoning scenarios. Overall, this work shows the effectiveness of diffusion-based modeling for logical reasoning, providing a unified and effective framework for knowledge graph reasoning.
However, the main limitation of our approach lies in its reliance on the ids of knowledge graph entities and relationships, lacking more flexible natural language question-answering capabilities. Furthermore, this method relies on a large number of hypothesis-observation pairs for sampling, as well as a long training time. This limits the model’s ability to transfer knowledge across different knowledge graphs and its scalability.
7. Acknowledgements
The corresponding author is Yangqiu Song. We owe sincerely thanks to all authors for their valuable efforts and contributions. The authors of this paper are supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China and the National Natural Science Foundation of China (No.62462007 and No.62302023).